# Comparative Analysis of Pooling Mechanisms in LLMs: A Sentiment Analysis Perspective

### v2

@@ Abstract
I Introduction
Large Language Models (LLMs) have emerged as transformative tools in natural language processing (NLP), offering unparalleled performance across a broad spectrum of tasks. Among these, BERT (Bidirectional Encoder Representations from Transformers) [1] and GPT (Generative Pre-trained Transformer) [2] stand out as two of the most influential architectures. BERT’s bidirectional attention mechanism enables it to deeply understand contextual relationships within text, making it particularly effective for comprehension-based tasks. On the other hand, GPT’s unidirectional design and autoregressive modeling excel in generating coherent and contextually appropriate text. Together, these models exemplify the state of the art in leveraging transformer-based architectures to tackle diverse linguistic challenges.
The applications of LLMs can be broadly categorized into token-level and sentence-level tasks [3]. Token-level tasks, such as named entity recognition and part-of-speech tagging, require the model to process and predict attributes for individual tokens within a sequence. Sentence-level tasks, including sentiment analysis, entail aggregating token-level information to derive an overall understanding of the text’s meaning. These tasks depend heavily on effective mechanisms to condense token embeddings into coherent sentence-level representations. In this study, we focus on sentence-level tasks, particularly sentiment analysis, as it serves as a fundamental benchmark for evaluating the semantic capabilities of LLMs.
Pooling layers play a critical role in sentence-level tasks by aggregating token embeddings into unified representations [4]. Commonly employed pooling strategies include Mean pooling, which averages embeddings to provide a balanced view, Max pooling, which captures the most salient features, and Weighted Sum pooling, which applies learned weights to emphasize contextually significant tokens. Despite their importance, the comparative effects of these pooling mechanisms on different LLM architectures remain underexplored. To bridge this gap, this paper evaluates the performance of these pooling strategies on BERT and GPT within the context of sentiment analysis. Our contributions are threefold:
@@ II Related Work
Deep Learning models, especially the Transformer-based [5, 6, 7] have found innovative applications across various fields like image analysis [8, 9], virtual reality [10], sequences modeling [11], and emotion recognition [12].
@@
The transformer architecture, proposed by Vaswani et al. [13], addressed the sequential computation limitations of RNNs through self-attention mechanisms. This innovation enabled parallel processing of input sequences and more effective modeling of long-range dependencies, establishing the foundation for modern LLMs. The transformer’s encoder-decoder architecture demonstrated superior performance in machine translation tasks and quickly became the de facto standard for neural sequence transduction models [14].
@@
BERT, introduced by Devlin et al. [1], marked a breakthrough in NLP by utilizing a bidirectional transformer architecture to generate contextual embeddings for tokens. Unlike unidirectional models, BERT captures relationships between words in both forward and backward directions, enabling superior understanding of text semantics. Numerous variations of BERT have since been developed to enhance its performance and efficiency. For example, RoBERTa [15] optimized BERT’s training process by removing the next sentence prediction task and training on a larger dataset. DistilBERT [16] focuses on reducing model size while retaining most of BERT’s performance, making it suitable for resource-constrained environments. BERT and its variants have been widely adopted for tasks such as sentiment analysis, text classification, and question answering.
@@
GPT, initially proposed by OpenAI [2], pioneered autoregressive modeling for text generation. The model predicts the next word in a sequence based on prior context, making it particularly effective for generative tasks such as text completion and summarization. Its successor, GPT-2 [17], significantly expanded the model’s capacity and demonstrated remarkable versatility across tasks without task-specific fine-tuning. The introduction of GPT-3 [18] and GPT-4 [19] further advanced the state of the art, leveraging billions of parameters to perform complex tasks such as language translation and dialogue generation with minimal instruction. GPT models have found extensive applications in content creation, conversational agents, and zero-shot learning scenarios [20].
@@
Pooling mechanisms are critical for aggregating token embeddings into sentence-level representations. Mean pooling, one of the simplest techniques, computes the average of token embeddings, providing a balanced view of the input. Max pooling, in contrast, captures the most salient features by selecting the maximum value along each dimension. Weighted Sum pooling introduces a learnable weighting mechanism to prioritize contextually important tokens [22, 21]. These pooling strategies have been explored in various contexts [21]. For instance, Justyna et al. [23] incorporated Mean pooling in their work on contextualized embeddings for sentence classification. Max pooling was applied by Conneau et al. [24] in supervised learning tasks to enhance feature selection. Despite these efforts, the comparative impact of pooling operations on LLMs such as BERT and GPT remains underexplored, especially for sentence-level tasks like sentiment analysis.
III Methodology
III-A Attention Mechanism Fundamentals
@@
The attention mechanism serves as the architectural cornerstone of transformer-based Large Language Models (LLMs) [4]. At its core, attention allows models to dynamically weight the importance of different tokens when processing sequential data. The standard attention mechanism can be mathematically formulated as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
@@
In practice, multi-head attention mechanisms [14], as shown in Figure 1, are employed to enhance model expressiveness and capture diverse token interactions. Multi-head attention involves projecting the input into multiple subspaces, each processed by an independent attention mechanism. The outputs are then concatenated and linearly transformed to produce the final attention output.
Figure 1: Multi-Head Attention Mechanism
III-B Pooling Strategies
@@
III-B1 Mean Pooling
$\text{Pool}_{\text{mean}}(X)=\frac{1}{l}\sum_{i=1}^{l}x_{i}$ (2)
@@ III-B1 Mean Pooling
III-B2 Max Pooling
$\text{Pool}_{\text{max}}(X)=\max_{i=1}^{l}x_{i}$ (3)
@@ III-B2 Max Pooling
III-B3 Weighted Sum Pooling
$\text{Pool}_{\text{ws}}(X)=\sum_{i=1}^{l}w_{i}x_{i}$ (4)
@@ IV-A Datasets
The deliberate equal distribution of positive and negative samples mitigates class imbalance, ensuring robust model training and evaluation. This balanced approach prevents potential biases that could arise from skewed class representations. Table I provides a summarization.
TABLE I: Dataset Summarization
@@ IV-B2 Evaluation Metrics
IV-C Results
IV-C1 Confusion Matrix Analysis
As shown in Figure 2-3, both BERT and GPT2 demonstrated comparable performance across pooling mechanisms, with an aggregate correct prediction rate of approximately 86.07
@@ IV-C1 Confusion Matrix Analysis
Figure 2: Confusion Matrix of Bert with Different Pooling Strategies
@@
Figure 3: Confusion Matrix of GPT2 with Different Pooling Strategies
@@
IV-C2 Performance Metrics Evaluation
@@
Figure 4: Precision, Recall, F1 Score of Bert and GPT2 with Different Pooling Strategies

### v3

@@ Abstract.
conference: April 26–28, 2025; Tokyo, Japan; ccs: Computing methodologies Machine learningccs: Information systems Data miningccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
1. Introduction
Large Language Models (LLMs) have emerged as transformative tools in natural language processing (NLP), offering unparalleled performance across a broad spectrum of tasks. Among these, BERT (Bidirectional Encoder Representations from Transformers) (bert18bert, ) and GPT (Generative Pre-trained Transformer) (gpt18improving, ) stand out as two of the most influential architectures. BERT’s bidirectional attention mechanism enables it to deeply understand contextual relationships within text, making it particularly effective for comprehension-based tasks. On the other hand, GPT’s unidirectional design and autoregressive modeling excel in generating coherent and contextually appropriate text. Together, these models exemplify the state of the art in leveraging transformer-based architectures to tackle diverse linguistic challenges.
The applications of LLMs can be broadly categorized into token-level and sentence-level tasks (floridi20gpt, ). Token-level tasks, such as named entity recognition and part-of-speech tagging, require the model to process and predict attributes for individual tokens within a sequence. Sentence-level tasks, including sentiment analysis, entail aggregating token-level information to derive an overall understanding of the text’s meaning. These tasks depend heavily on effective mechanisms to condense token embeddings into coherent sentence-level representations. In this study, we focus on sentence-level tasks, particularly sentiment analysis, as it serves as a fundamental benchmark for evaluating the semantic capabilities of LLMs.
Pooling layers play a critical role in sentence-level tasks by aggregating token embeddings into unified representations (yao24survey, ). Commonly employed pooling strategies include Mean pooling, which averages embeddings to provide a balanced view, Max pooling, which captures the most salient features, and Weighted Sum pooling, which applies learned weights to emphasize contextually significant tokens. Despite their importance, the comparative effects of these pooling mechanisms on different LLM architectures remain underexplored. To bridge this gap, this paper evaluates the performance of these pooling strategies on BERT and GPT within the context of sentiment analysis. Our contributions are threefold:
@@ 2. Related Work
Deep Learning models (cheng24patch, ; cheng25unifying, ; cheng22deep, ; cheng22estimation, ; xing24traffic, ; cheng23shapnn, ), especially the Transformer-based (yangbotnet, ; linresearch, ; yangrac, ; li24contextual, ; luo24rmgen, ) have found innovative applications across various fields like image analysis (zhaoresearch, ; yangcnn, ), virtual reality (yangzhan, ), sequences modeling (dengresearch, ), medical diagnosis (yang24tcell, ), and emotion recognition (yangarxiv1, ).
@@
The transformer architecture, proposed by Vaswani et al. (vas17attention, ), addressed the sequential computation limitations of RNNs through self-attention mechanisms. This innovation enabled parallel processing of input sequences and more effective modeling of long-range dependencies, establishing the foundation for modern LLMs. The transformer’s encoder-decoder architecture demonstrated superior performance in machine translation tasks and quickly became the de facto standard for neural sequence transduction models (xing24enhancing, ).
@@
BERT, introduced by Devlin et al. (bert18bert, ), marked a breakthrough in NLP by utilizing a bidirectional transformer architecture to generate contextual embeddings for tokens. Unlike unidirectional models, BERT captures relationships between words in both forward and backward directions, enabling superior understanding of text semantics. Numerous variations of BERT have since been developed to enhance its performance and efficiency. For example, RoBERTa (liu19roberta, ) optimized BERT’s training process by removing the next sentence prediction task and training on a larger dataset. DistilBERT (sanh19distil, ) focuses on reducing model size while retaining most of BERT’s performance, making it suitable for resource-constrained environments. BERT and its variants have been widely adopted for tasks such as sentiment analysis, text classification, and question answering.
@@
GPT, initially proposed by OpenAI (gpt18improving, ), pioneered autoregressive modeling for text generation. The model predicts the next word in a sequence based on prior context, making it particularly effective for generative tasks such as text completion and summarization. Its successor, GPT-2 (radford19language, ), significantly expanded the model’s capacity and demonstrated remarkable versatility across tasks without task-specific fine-tuning. The introduction of GPT-3 (brown20language, ) and GPT-4 (achiam23gpt, ) further advanced the state of the art, leveraging billions of parameters to perform complex tasks such as language translation and dialogue generation with minimal instruction. GPT models have found extensive applications in content creation, conversational agents, and zero-shot learning scenarios (lund23chatting, ).
@@
Pooling mechanisms are critical for aggregating token embeddings into sentence-level representations. Mean pooling, one of the simplest techniques, computes the average of token embeddings, providing a balanced view of the input. Max pooling, in contrast, captures the most salient features by selecting the maximum value along each dimension. Weighted Sum pooling introduces a learnable weighting mechanism to prioritize contextually important tokens (gao22param, ; xing22weighted, ). These pooling strategies have been explored in various contexts (xing22weighted, ). For instance, Justyna et al. (sar21detection, ) incorporated Mean pooling in their work on contextualized embeddings for sentence classification. Max pooling was applied by Conneau et al. (conneau17supervised, ) in supervised learning tasks to enhance feature selection. Despite these efforts, the comparative impact of pooling operations on LLMs such as BERT and GPT remains underexplored, especially for sentence-level tasks like sentiment analysis.
3. Methodology
3.1. Attention Mechanism Fundamentals
@@
The attention mechanism serves as the architectural cornerstone of transformer-based Large Language Models (LLMs) (yao24survey, ). At its core, attention allows models to dynamically weight the importance of different tokens when processing sequential data. The standard attention mechanism can be mathematically formulated as:
(1) $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
@@
In practice, multi-head attention mechanisms (xing24enhancing, ), as shown in Figure 1, are employed to enhance model expressiveness and capture diverse token interactions. Multi-head attention involves projecting the input into multiple subspaces, each processed by an independent attention mechanism. The outputs are then concatenated and linearly transformed to produce the final attention output.
Figure 1. Multi-Head Attention Mechanism
3.2. Pooling Strategies
@@
3.2.1. Mean Pooling
(2) $\text{Pool}_{\text{mean}}(X)=\frac{1}{l}\sum_{i=1}^{l}x_{i}$
@@ 3.2.1. Mean Pooling
3.2.2. Max Pooling
(3) $\text{Pool}_{\text{max}}(X)=\max_{i=1}^{l}x_{i}$
@@ 3.2.2. Max Pooling
3.2.3. Weighted Sum Pooling
(4) $\text{Pool}_{\text{ws}}(X)=\sum_{i=1}^{l}w_{i}x_{i}$
@@ 4.1. Datasets
The deliberate equal distribution of positive and negative samples mitigates class imbalance, ensuring robust model training and evaluation. This balanced approach prevents potential biases that could arise from skewed class representations. Table 1 provides a summarization.
Table 1. Dataset Summarization
@@ 4.2.2. Evaluation Metrics
4.3. Results
4.3.1. Confusion Matrix Analysis
As shown in Figure 2-3, both BERT and GPT2 demonstrated comparable performance across pooling mechanisms, with an aggregate correct prediction rate of approximately 86.07%.
@@ 4.3.1. Confusion Matrix Analysis
Figure 2. Confusion Matrix of Bert with Different Pooling Strategies
@@
Figure 3. Confusion Matrix of GPT2 with Different Pooling Strategies
@@
4.3.2. Performance Metrics Evaluation
@@
Figure 4. Precision, Recall, F1 Score of Bert and GPT2 with Different Pooling Strategies